40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 11 12:50:10 2009, Jenne, Update, PEM, Seismometer Noise Characteristics HuddleTest_Aug2009data.png
    Reply  Thu Apr 22 20:28:40 2010, Jenne, Update, PEM, Seismometer Noise Characteristics Gur1_Gur2_noise.png
       Reply  Thu May 27 23:59:54 2010, rana, Update, PEM, New Foam Box installed Untitled.png
          Reply  Fri May 28 00:40:53 2010, rana, Update, PEM, DAQ down 
             Reply  Fri May 28 10:44:47 2010, josephb, Update, PEM, DAQ down 
                Reply  Tue Jun 8 18:39:36 2010, rana, Update, PEM, DAQ down 
                   Reply  Tue Jun 8 20:52:25 2010, josephb, Update, PEM, DAQ up (for the moment) 
          Reply  Fri May 28 11:26:35 2010, Jenne, Update, PEM, New Foam Box installed 
          Reply  Tue Jun 1 11:47:14 2010, steve, Update, PEM, lead balls on concrete P1060269.JPG
             Reply  Wed Jun 9 19:47:08 2010, nancy, Update, PEM, lead balls on concrete 
                Reply  Thu Jun 10 07:53:14 2010, Alberto, Update, PEM, LaTeXlabs 
                   Reply  Thu Jun 10 10:58:02 2010, Koji, Update, General, LaTeXlabs 
                      Reply  Thu Jun 10 11:10:21 2010, Alberto, Update, General, LaTeXlabs 
Message ID: 2237     Entry time: Wed Nov 11 12:50:10 2009     Reply to this: 2833
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Seismometer Noise Characteristics 

The attached plot shows the spectra of the 3 Z axes of the 3 seismometers we have (this data is from ~20Aug2009, when the Ranger was in the Z orientation) in Magenta, Cyan and Green, and the noise of each of the sensors in Red, Blue and Black.  The noise curves were extracted from the spectra using the Huddle Test / 3 Corner Hat method.  The Blue and Black traces which are just a few points are estimates of the noise from other spectra.  The Blue points come from the Guralp Spec Sheet, and the Black comes from the noise test that Rana and I did the other day with the Ranger (elog 2223).  

I'm not really happy with the black spectra - it looks way too high.  I'm still investigating to see if this is a problem with my calibration/method....

Attachment 1: HuddleTest_Aug2009data.png  32 kB  | Hide | Hide all
HuddleTest_Aug2009data.png
ELOG V3.1.3-